OpenStudy (ledah):

Solve for n: 8(n – 3) + 4n = 12n + 3

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

o..no solution

OpenStudy (jackellyn):

Distribute 8(n) - 8(3) + 4n = 12n + 3 8n -24 +4n = 12n +3 12n -24 = 12n +3 12n -12n -24 +24 = 12n -12n +3 +24 \[0 \neq 27\] Therefore, no solution.